OpenStudy (anonymous):

find lim as x approaches infinite of 5+radicalx/ 3x+1 if it exist

OpenStudy (anonymous):

\[\lim_{x\to \infty}\frac{5+\sqrt{x}}{3x+1}\]?

OpenStudy (anonymous):

if that is it, the answer is zero, because the numerator has degree \(\frac{1}{2}\) and the denominator has degree \(1\) imagine \(x=1,000,000\) the numerator would be \(1005\) and the denominator would be \(3000001\)

OpenStudy (anonymous):

see i was confused by the part '' if it exists" i thought they wanted me to check from right and left of infinity

OpenStudy (anonymous):

there is only one way to get to positive infinity

OpenStudy (anonymous):

Thanks a lot

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

use continuity to evaluate lim as x approaches pie of sin(x+ sinx)

OpenStudy (anonymous):

since this is a continuous function (composition of continuous functions is continuous) you job it only to replace \(x\) by \(\pi\) and compute \[\sin(\pi+\sin(\pi))\]

OpenStudy (anonymous):

Wow thanks once again

OpenStudy (anonymous):

yw